& A 10 Ω resistor, a 15 Ω resistor, and a 5 Ω resistor are connected in series across a 90 V battery.

a. What is the equivalent resistance of the circuit?

b. What is the current in the circuit?
1
Expert's answer
2012-03-20T11:08:27-0400
R=R1+R2+R3=30
I=U/R=30A